A stone after having fallen from rest under the influence of gravity for 6 seconds crashes throught a horizontal glass plate, thereby losing 2/3rd of its velocity. If it  then reaches the ground in 2 seconds, find the height of the plate above the ground?

Grade:11

2 Answers

Khimraj
5 years ago
velocity of stone after 6 sec = gt = 10*6 = 60 m/s
velocity AFTER CRASHING WITH GLASS = (1/3)*60 = 20m/s
now stone reaches ground in 2 sec
so
h = 20*2 + ½*10*22 = 40 + 20 = 60m.
Hope it clears.
Rajdeep
5 years ago
HELLO!
 
Given, the stone starts from rest (for the free fall).
Hence, u = 0.
Time for which it falls, t = 6 seconds.
The final velocity v, just before striking the glass plate, needs to be found out.
 
From relation v = u + at,
Putting u = 0 and a = g = 10m/s2
v = gt = 10 x 6 = 60m/s
 
Now, after striking, it loses 2/3rd of its velocity.
 
So, velocity it loses is 2/3 x 60 m/s = 40 m/s
 
Now, velocity = 60 – 40 = 20 m/s
 
For the second part of the motion,
take u = 20 m/s
Time t = 2 seconds.
 
Height of plate from ground,
h = ut + ½ gt2
 
or h = 20 x 2 + ½ x 10 x 4
= 40 + 20
= 60m
 
Hence, height of the glass place from the ground is 60 metres.